### v3

# Null Distance and Temporal Functions

Abstract
The notion of null distance was introduced by Sormani and Vega in [1] as part of a broader program to develop a theory of metric convergence adapted to Lorentzian geometry. Given a time function $\tau$ on a spacetime $(M,g)$, the associated null distance $\hat{d}_{\tau}$ is constructed from and closely related to the causal structure of $M$. While generally only a semi-metric, $\hat{d}_{\tau}$ becomes a metric when $\tau$ satisfies the local anti-Lipschitz condition.
In this work, we focus on temporal functions, that is, differentiable functions whose gradient is everywhere past-directed timelike. Sormani and Vega showed that the class of $C^{1}$ temporal functions coincides with that of $C^{1}$ locally anti-Lipschitz time functions. When a temporal function $f$ is smooth, its level sets $M_{t}=f^{-1}(t)$ are spacelike hypersurfaces and thus Riemannian manifolds endowed with the induced metric $h_{t}$. Our main result establishes that, on any level set $M_{t}$ where the gradient $\nabla f$ has constant norm, the null distance $\hat{d}_{f}$ is bounded above by a constant multiple of the Riemannian distance $d_{h_{t}}$.
Applying this result to a smooth regular cosmological time function $\tau_{g}$—as introduced by Andersson, Galloway, and Howard in [2]—we prove a theorem confirming a conjecture of Sakovich and Sormani [3]: if the diameters of the level sets $M_{t}=\tau_{g}^{-1}(t)$ shrink to zero as $t\to 0$, then the spacetime exhibits a Big Bang singularity, as defined in their work.
1 Introduction
Let $(M,g)$ be a spacetime, that is, a connected, time-oriented Lorentzian manifold. A continuous function $\tau\colon M\to\mathbb{R}$ is called a time function if it is strictly increasing along all future-directed causal curves. The null distance associated with a time function was defined by Sormani and Vega as follows. [^1] In [1], the null distance is defined with respect to a generalized time function, which is not necessarily continuous. In this work, we restrict our attention to continuous time functions.
Definition 1.1.
A curve $\beta\colon[a,b]\to M$ is called a piecewise causal curve if there exists a partition $a=s_{0}<\dots<s_{N}=b$ such that each restriction $\beta_{i}={\beta}|_{[s_{i},s_{i+1}]}$ is either a future-directed or a past-directed causal curve. Given a time function $\tau$ on $M$, the null length of a piecewise causal curve $\beta$ is defined by
$\hat{L}_{\tau}(\beta)=\sum_{i=0}^{N-1}\left|{\tau(\beta(s_{i+1}))-\tau(\beta(s_{i}))}\right|.$
The null distance between two points $p,q\in M$ is then given by
[…]
If $q=p$, we set $\hat{d}_{\tau}(p,q)=0$.
It follows readily that $\hat{d}_{\tau}$ is a semi-metric on $M$. However, in general, it fails to be a metric, as it may vanish for distinct points. Sormani and Vega identified a condition on the time function—the local anti-Lipschitz condition—which is equivalent to the definiteness of $\hat{d}_{\tau}$. When $\tau$ satisfies this condition, the associated null distance $\hat{d}_{\tau}$ is a metric on $M$ that induces the manifold topology.
A notable example of a time function is a temporal function, that is, a differentiable function $f\colon M\to\mathbb{R}$ whose gradient $\nabla f$ is everywhere past-directed timelike. [^2] We slightly deviate from the standard definition of a temporal function, which requires smoothness. Sormani and Vega proved that the class of $C^{1}$ temporal functions coincides with that of $C^{1}$ locally anti-Lipschitz time functions. Consequently, if $f$ is a $C^{1}$ temporal function, then the associated null distance $\hat{d}_{f}$ is a metric on $M$ that induces the manifold topology. Furthermore, if $f$ is smooth and the level set $M_{t}=f^{-1}(t)$ is nonempty for some $t\in\mathbb{R}$, then $M_{t}$ is a spacelike hypersurface. Hence, the restriction of the ambient Lorentzian metric $g$ to $M_{t}$, denoted by $h_{t}$, is a Riemannian metric on $M_{t}$. The following theorem represents our main result.
Theorem 1.2.
Let $(M,g)$ be a spacetime admitting a smooth temporal function $f\colon M\to\mathbb{R}$. Suppose there exist $t\in\mathbb{R}$ and $C>0$ such that the level set $M_{t}=f^{-1}(t)$ is nonempty and the equality
$g(\nabla f,\nabla f)=-C^{2}$
holds at every point of $M_{t}$. Then, for every $p,q\in M_{t}$,
$\hat{d}_{f}(p,q)\leq Cd_{h_{t}}(p,q),$
where $h_{t}$ denotes the induced Riemannian metric on $M_{t}$ and $d_{h_{t}}$ is the corresponding distance function.
An example of a smooth temporal function with unit gradient everywhere is a smooth regular cosmological time function $\tau_{g}\colon M\to(0,\infty)$, as introduced by Andersson, Galloway, and Howard in [2]. Like any locally anti-Lipschitz time function, $\tau_{g}$ is $1$-Lipschitz with respect to the associated null distance $\hat{d}_{\tau_{g}}$, and hence admits a unique $1$-Lipschitz extension $\overline{\tau_{g}}\colon\overline{M}\to[0,\infty)$ to the metric completion of $(M,\hat{d}_{\tau_{g}})$. This extension allows the definition of the initial level set $\overline{\tau_{g}}^{-1}(0)\subseteq\overline{M}$.
In [3], Sakovich and Sormani considered the case in which the initial level set consists of a single point, in connection with their proposal for a general class of Big Bang spacetimes. Building upon Theorem 1.2, we establish the following result, which confirms Conjecture 3.7 of [3].
Theorem 1.3.
Let $(M,g)$ be a spacetime admitting a smooth regular cosmological time function $\tau_{g}\colon M\to(0,\infty)$. For each $t>0$, if the level set $M_{t}=\tau_{g}^{-1}(t)$ is nonempty, denote by $h_{t}$ the induced Riemannian metric on $M_{t}$ and by $d_{h_{t}}$ the corresponding distance function. If
$\lim_{t\to 0^{+}}\operatorname{diam}_{d_{h_{t}}}(M_{t})=0,$
then the initial level set $\overline{\tau_{g}}^{-1}(0)\subseteq\overline{M}$ consists of a single point $p_{BB}\in\overline{M}$.
Outline.
Section 2 introduces definitions and results from Lorentzian and metric geometry that will be used throughout. Section 3 focuses on the notion of null distance. Section 4 presents the proof of Theorem 1.2. Finally, Section 5 discusses regular cosmological time functions and establishes Theorem 1.3.
Acknowledgments.
Part of this work was carried out during a research stay at the University of Vienna in the group of Prof. Michael Kunzinger, which is partly supported by the Austrian Science Fund (FWF) [Grant DOI 10.55776/EFP6]. I am grateful to Chiara Rigoni for giving me the opportunity to join the group and for proposing the topic. I am especially thankful to Tobias Beran, whose insights and ideas have been a valuable source of inspiration. I also wish to thank Professors Christina Sormani and Anna Sakovich for their interest in this work. Finally, I am sincerely grateful to my thesis advisor, Prof. Stefano Pigola, for welcoming my research proposal with openness and for engaging in constructive discussions that contributed to its improvement.
2 Preliminaries
2.1 Lorentzian Background
This subsection provides a brief overview of Lorentzian geometry, presenting definitions and results used throughout. For a comprehensive treatment of the basic theory, see [4].
2.1.1 Spacetimes
A Lorentzian manifold is a pair $(M,g)$, where $M$ is a smooth manifold of dimension $n+1$ and $g$ is a Lorentzian metric on $M$, that is, a metric tensor of signature $(1,n)$. Each non-zero tangent vector $X\in TM$ is classified as timelike, null, or spacelike, according to whether $g(X,X)$ is negative, zero, or positive, respectively. Timelike and null vectors are collectively referred to as causal. The $g$-norm of a tangent vector $X\in TM$ is defined by
$\lVert{X}\rVert_{g}=\sqrt{\left|{g(X,X)}\right|}.$
Causal vectors satisfy a reverse form of the Cauchy–Schwarz inequality.
Proposition 2.1 (Reverse Cauchy–Schwarz Inequality).
Let $X,Y\in TM$ be causal vectors at the same point. Then
$g(X,Y)^{2}\geq g(X,X)g(Y,Y),$
with equality if and only if $X$ and $Y$ are linearly dependent.
A time orientation on a Lorentzian manifold $(M,g)$ is a smooth timelike vector field $\vartheta$ on $M$. A causal vector $X\in TM$ is future-directed (respectively, past-directed) if $g(X,\vartheta)<0$ (respectively, $g(X,\vartheta)>0$). The following lemma ensures that every causal vector must be either future-directed or past-directed. The proof is provided in Appendix 6.1.
Lemma 2.2.
Let $X,T\in TM$ be tangent vectors at the same point, with $X$ causal and $T$ timelike. Then $g(X,T)\neq 0$.
The following characterization of past-directed timelike vectors will be used throughout this work. The proof is provided in Appendix 6.1.
Lemma 2.3.
Let $p\in M$, and let $h_{p}$ be any Euclidean inner product on $T_{p}M$. Then the following conditions on a tangent vector $T\in T_{p}M$ are equivalent:
  1. (1)
    $T$ is past-directed timelike.
  2. (2)
    For every future-directed causal vector $X\in T_{p}M$, $g(X,T)>0$.
  3. (3)
    There exists a constant $C>0$ such that, for all future-directed causal vectors $X\in T_{p}M$,
    $g(T,X)\geq C\lVert{X}\rVert_{h_{p}}.$
A spacetime is a connected, time-oriented Lorentzian manifold. Unless otherwise specified, the pair $(M,g)$ will denote a spacetime.
Example 2.4.
The most basic example of a spacetime is the $(n+1)$-dimensional Minkowski spacetime:
$\mathbb{M}^{n+1}=(\mathbb{R}^{n+1},\eta),$
where
$\eta=-dt^{2}+d(x^{1})^{2}+\dots+d(x^{n})^{2}$
is the Minkowski metric expressed in Cartesian coordinates. The time orientation is given by the coordinate vector field $\partial_{t}$.
Throughout this work, we denote points in $\mathbb{M}^{n+1}$ by $p=(t_{p},x_{p})$, where $t_{p}\in\mathbb{R}$ is the temporal component and $x_{p}\in\mathbb{R}^{n}$ the spatial one. Moreover, given two points $p,q\in\mathbb{M}^{n+1}$, we denote by $[p,q]$ the straight-line segment from $p$ to $q$, regarded as a parametrized curve.
Example 2.5.
A generalized Robertson–Walker (GRW) spacetime is a spacetime of the form
$(M,g)=(I\times S,-dt^{2}+f(t)^{2}h),$
where $I\subseteq\mathbb{R}$ is an open interval, $(S,h)$ is an $n$-dimensional connected Riemannian manifold, and $f\colon I\to\mathbb{R}$ is a smooth, strictly positive function called the scale factor. The time orientation is given by the coordinate vector field $\partial_{t}$.
Note that Minkowski spacetime $\mathbb{M}^{n+1}$ arises as a special case of a GRW spacetime by taking $I=\mathbb{R}$, $f\equiv 1$, and $(S,h)$ equal to the $n$-dimensional Euclidean space.
2.1.2 Causality
A smooth curve $\gamma\colon I\to M$ is called timelike (respectively, null, spacelike, causal) if its tangent vector $\gamma^{\prime}$ is everywhere timelike (respectively, null, spacelike, causal). A causal curve is future-directed (respectively, past-directed) if its tangent vector is everywhere future-directed (respectively, past-directed).
The causal future of a point $p\in M$ is defined by
[…]
Similarly, the chronological future of $p$ is given by
[…]
The causal past $J^{-}(p)$ and the chronological past $I^{-}(p)$ are defined analogously, replacing future-directed curves with past-directed ones. For any $p,q\in M$, we write $p\leq q$ if and only if $q\in J^{+}(p)$, or equivalently, $p\in J^{-}(q)$.
Proposition 2.6.
For every point $p\in M$, the chronological future $I^{+}(p)$ and the chronological past $I^{-}(p)$ are open subsets of $M$.
2.1.3 Time Functions
A continuous function $\tau\colon M\to\mathbb{R}$ is called a time function if it is strictly increasing along all future-directed causal curves.
Recall that, if a function $f\colon M\to\mathbb{R}$ is differentiable at a point $p\in M$, then its gradient at $p$ is the unique tangent vector satisfying
[…]
The following lemma shows that if a time function is differentiable at a point, then its monotonicity along future-directed causal curves imposes constraints on the causal character of its gradient. The proof is provided in Appendix 6.1.
Lemma 2.7.
Let $\tau$ be a time function on $M$. If $\tau$ is differentiable at a point $p\in M$, then the gradient vector $(\nabla\tau)_{p}$ is either past-directed causal or zero.
As a consequence, the gradient of a differentiable time function is either past-directed causal or zero at every point. However, it is not necessarily timelike everywhere. A differentiable function whose gradient is everywhere past-directed timelike is called a temporal function. Every temporal function is, in particular, a time function.
Proposition 2.8.
Let $f\colon M\to\mathbb{R}$ be a temporal function. Then $f$ is a time function.
Proof.
Let $\alpha$ be a future-directed causal curve. Since $\nabla f$ is past-directed timelike and $\alpha^{\prime}$ is future-directed casual, Lemma 2.3 implies that
$(f\circ\alpha)^{\prime}=g(\nabla f,\alpha^{\prime})>0.$
It follows that $f$ is strictly increasing along $\alpha$, and thus a time function. ∎
Example 2.9.
In a GRW spacetime
$(M,g)=(I\times S,-dt^{2}+f(t)^{2}h),$
consider a differentiable function of the form $\tau(t,x)=\phi(t)$, where $\phi\colon I\to\mathbb{R}$ satisfies $\phi^{\prime}(t)>0$ for all $t\in I$. Then $\tau$ is a temporal function. Indeed, its gradient is given by
$\nabla\tau=-\phi^{\prime}(t)\partial_{t},$
which is everywhere past-directed timelike.
We now briefly recall some standard causality conditions. A spacetime $(M,g)$ is said to be causal if it contains no closed causal curves, and stably causal if this property persists under small perturbations of the metric $g$. Every stably causal spacetime is, in particular, strongly causal, meaning that for every point $p\in M$ and every open neighborhood $U$ of $p$, there exists an open neighborhood $V\subseteq U$ of $p$ such that every causal curve with endpoints in $V$ is entirely contained in $U$. The following is a fundamental result in Lorentzian geometry.
Theorem 2.10.
Let $(M,g)$ be a spacetime. Then the following conditions are equivalent:
  1. (1)
    $(M,g)$ admits a time function.
  2. (2)
    $(M,g)$ admits a smooth temporal function.
  3. (3)
    $(M,g)$ is stably causal.
Proof.
See [5, Thm. 4.12]. ∎
2.1.4 Lorentzian Distance
The Lorentzian length of a causal curve $\gamma\colon I\to M$ is defined by
$L_{g}(\gamma)=\int_{I}\sqrt{-g(\gamma^{\prime}(s),\gamma^{\prime}(s))}\,ds.$
The Lorentzian distance between two points $p,q\in M$ is then given by
[…]
If $q=p$ or $q\notin J^{+}(p)$, we set $d_{g}(p,q)=0$. Despite its name, the Lorentzian distance is not a metric, as it may vanish for distinct points and is generally not symmetric, except in trivial cases.
2.2 Metric Background
This subsection introduces definitions and results from metric geometry that will be used throughout. We adopt much of the notation and terminology of [6], to which we also refer for many of the proofs.
2.2.1 Metric Spaces
Definition 2.11.
Let $X$ be a set. A function $d\colon X\times X\to\mathbb{R}\cup\{\infty\}$ is called a metric on $X$ if the following conditions hold for all $x,y,z\in X$:
  1. (1)
    Positiveness: $d(x,y)\geq 0$, with equality if and only if $x=y$.
  2. (2)
    Symmetry: $d(x,y)=d(y,x)$.
  3. (3)
    Triangle inequality: $d(x,z)\leq d(x,y)+d(y,z)$.
The condition that $d(x,y)=0$ implies $x=y$ is referred to as definiteness. If $d$ satisfies all the conditions above except definiteness, it is called a semi-metric on $X$.
A metric space is a pair $(X,d)$, where $X$ is a set and $d$ is a metric on it. Unless multiple metrics on the same set $X$ are considered, we will omit an explicit reference to the metric and write "a metric space XX" instead of "a metric space (X,d)(X,d)".
For any $x\in X$ and $r>0$, we denote by $B_{r}(x)$ the open ball of radius $r$ centered at $x$, that is, the set of points $y\in X$ such that $d(x,y)<r$. Given a nonempty subset $S\subseteq X$, its diameter is defined by
$\operatorname{diam}(S)=\sup\{d(x,y)\mid x,y\in S\}.$
If $\operatorname{diam}(S)<\infty$, the set $S$ is said to be bounded. Given two nonempty subsets $A,B\subseteq X$, their distance is defined by
$\operatorname{dist}(A,B)=\inf\{d(a,b)\mid a\in A,\ b\in B\}.$
The following result, known as Lebesgue’s number lemma, will be used throughout this work.
Proposition 2.12 (Lebesgue’s Number Lemma).
Let $X$ be a compact metric space, and let $\{U_{\alpha}\}_{\alpha\in A}$ be an open cover of $X$. Then there exists $\varepsilon>0$ such that every ball of radius $\varepsilon$ in $X$ is entirely contained in some $U_{\alpha}$.
Proof.
See [6, Thm. 1.6.11]. ∎
2.2.2 Hausdorff Distance
For any nonempty $S\subseteq X$ and $r>0$, we denote by $U_{r}(S)$ the tubular open neighborhood of radius $r$ around $S$, that is, the set of points $x\in X$ such that $\operatorname{dist}(x,S)<r$. Equivalently,
$U_{r}(S)=\bigcup_{x\in S}B_{r}(x).$
Given two nonempty subsets $A,B\subseteq X$, their Hausdorff distance is defined by
$d_{H}(A,B)=\inf\{r>0\mid A\subseteq U_{r}(B),\ B\subseteq U_{r}(A)\}.$
It is straightforward to verify the following equivalent expression:
$d_{H}(A,B)=\max\{\sup\nolimits_{a\in A}\operatorname{dist}(a,B),\ \sup\nolimits_{b\in B}\operatorname{dist}(b,A)\}.$
Proposition 2.13.
Let $X$ be a metric space. Then the Hausdorff distance is a semi-metric on the set of all nonempty subsets of $X$. Moreover, if $A$ and $B$ are closed nonempty subsets of $X$ such that $d_{H}(A,B)=0$, then $A=B$.
Proof.
See [6, Prop. 7.3.3]. ∎
We denote by $\mathcal{F}(X)$ the set of all nonempty closed subsets of $X$ (from the French fermé), and by $\mathcal{F}_{b}(X)$ the set of all nonempty closed and bounded subsets of $X$. The previous proposition shows that the Hausdorff distance is a metric on $\mathcal{F}(X)$. Moreover, it is straightforward to verify that $\mathcal{F}_{b}(X)$ is a closed subset of $\mathcal{F}(X)$.
Lemma 2.14.
Let $X$ be a metric space, and let $\operatorname{diam}\colon\mathcal{F}_{b}(X)\to[0,\infty)$ denote the function that assigns to each set its diameter. Then $\operatorname{diam}$ is $2$-Lipschitz with respect to the Hausdorff distance.
Proof.
Let $A,B\in\mathcal{F}_{b}(X)$, and set $d=d_{H}(A,B)$. By the definition of the Hausdorff distance, for every $\varepsilon>0$ we have
$A\subseteq U_{d+\varepsilon}(B)\quad\text{and}\quad B\subseteq U_{d+\varepsilon}(A).$
This implies
$\left|{\operatorname{diam}(A)-\operatorname{diam}(B)}\right|\leq 2(d_{H}(A,B)+\varepsilon).$
The conclusion follows by letting $\varepsilon\to 0$. ∎
The following lemma follows directly from the continuity of $\operatorname{diam}$ with respect to the Hausdorff distance and the fact that $\mathcal{F}_{b}(X)$ is closed in $\mathcal{F}(X)$.
Lemma 2.15.
Let $X$ be a metric space, and let $A_{j}\in\mathcal{F}_{b}(X)$ be a sequence of nonempty closed and bounded subsets. If $A_{j}$ converges to a set $A\in\mathcal{F}(X)$ in the Hausdorff sense, then $A\in\mathcal{F}_{b}(X)$ and
$\operatorname{diam}(A)=\lim_{j\to\infty}\operatorname{diam}(A_{j}).$
2.2.3 Metric Completion
Proposition 2.16.
Let $X$ be a metric space. Then there exist a complete metric space $\overline{X}$ and an isometric embedding of $X$ as a dense subset of $\overline{X}$. Moreover, any other complete metric space containing an isometric copy of $X$ as a dense subset is isometric to $\overline{X}$.
Proof.
See [6, Thm. 1.5.10]. ∎
The space $\overline{X}$ is called the metric completion of $X$. In what follows, we identify $X$ with its isometric copy in $\overline{X}$, writing $X\subseteq\overline{X}$ and using the same symbol $d$ to denote both the original metric on $X$ and its extension to $\overline{X}$.
Proposition 2.17.
Let $X$ be a metric space, and let $S\subseteq X$ be a dense subset. Let $Y$ be a complete metric space, and let $f\colon S\to Y$ be a $C$-Lipschitz function for some $C>0$. Then there exists a unique $C$-Lipschitz function $\tilde{f}\colon X\to Y$ such that ${\tilde{f}}|_{S}=f$.
Proof.
See [6, Prop. 1.5.9]. ∎
Corollary 2.18.
Let $f\colon X\to Y$ be a $C$-Lipschitz function from a metric space $X$ to a complete metric space $Y$. Then $f$ admits a unique $C$-Lipschitz extension $\overline{f}\colon\overline{X}\to Y$ to the metric completion of $X$.
3 Null Distance
This section focuses on the notion of null distance, as introduced by Sormani and Vega in [1]. Definitions, key results, and proofs are drawn from their work, sometimes adapted for the present context.
3.1 Definitions
Definition 3.1.
A curve $\beta\colon[a,b]\to M$ is called a piecewise causal curve if there exists a partition $a=s_{0}<\dots<s_{N}=b$ such that each restriction $\beta_{i}={\beta}|_{[s_{i},s_{i+1}]}$ is either a future-directed or a past-directed causal curve.
Definition 3.2.
Let $\tau$ be a time function on $M$. The null length of a piecewise causal curve $\beta\colon[a,b]\to M$, with breakpoints $a=s_{0}<\dots<s_{N}=b$, is defined by
$\hat{L}_{\tau}(\beta)=\sum_{i=0}^{N-1}\left|{\tau(\beta(s_{i+1}))-\tau(\beta(s_{i}))}\right|.$
The null distance between two points $p,q\in M$ is then given by
[…]
If $q=p$, we set $\hat{d}_{\tau}(p,q)=0$.
The following lemma ensures that the null distance between any two points is well-defined.
Lemma 3.3.
For any pair of points $p,q\in M$, there exists a piecewise causal curve from $p$ to $q$.
Proof.
Let $\gamma\colon[a,b]\to M$ be a smooth curve with $\gamma(a)=p$ and $\gamma(b)=q$. Since
$M=\bigcup_{x\in M}I^{-}(x),$
and each set $I^{-}(x)$ is open by Proposition 2.6, the collection $\{\gamma^{-1}(I^{-}(x))\}_{x\in M}$ forms an open cover of the compact interval $[a,b]$. By the Lebesgue’s number lemma, there exists $\varepsilon>0$ such that every subinterval of $[a,b]$ of length less than $\varepsilon$ is contained in some $\gamma^{-1}(I^{-}(x))$.
Let $a\leq t_{0}<\dots<t_{N-1}\leq b$ be a finite collection of points such that the open intervals $(t_{i}-\varepsilon/2,\ t_{i}+\varepsilon/2)$ cover $[a,b]$. For each $i=1,\dots,N-1$, choose a point $s_{i}\in[a,b]$ such that
$t_{i-1}<s_{i}<t_{i},\quad s_{i}-t_{i-1}<\frac{\varepsilon}{2},\quad\text{and}\quad t_{i}-s_{i}<\frac{\varepsilon}{2}.$
Define $s_{0}=a$ and $s_{N}=b$. Then each interval $s_{i},s_{i+1}]$ has length less than $\varepsilon$, and is therefore contained in $\gamma^{-1}(I^{-}(x_{i}))$ for some $x_{i}\in M$. It follows that both endpoints $\gamma(s_{i})$ and $\gamma(s_{i+1})$ lie in $I^{-}(x_{i})$. Consequently, there exist a future-directed causal curve from $\gamma(s_{i})$ to $x_{i}$ and a past-directed causal curve from $x_{i}$ to $\gamma(s_{i+1})$. Their concatenation yields a piecewise causal curve from $\gamma(s_{i})$ to $\gamma(s_{i+1})$, with a breakpoint at $x_{i}$. The full concatenation of these segments is a piecewise causal curve from $\gamma(s_{0})=p$ to $\gamma(s_{N})=q$. ∎
3.2 Basic Properties
Lemma 3.4.
Let $\tau$ be a time function on $M$, and let $\beta\colon[a,b]\to M$ be a piecewise causal curve from $p$ to $q$, with breakpoints $a=s_{0}<\dots<s_{N}=b$. Then:
  1. (1)
    $\hat{L}_{\tau}(\beta)=\tau(q)-\tau(p)$ if and only if $\beta$ is future-directed causal.
  2. (2)
    The following inequalities hold:
    $\hat{L}_{\tau}(\beta)\geq\max_{s\in[a,b]}\tau(\beta(s))-\min_{s\in[a,b]}\tau(\beta(s))\geq\left|{\tau(q)-\tau(p)}\right|.$
  3. (3)
    If $\tau$ is $C^{1}$ along $\beta$, then
    $\hat{L}_{\tau}(\beta)=\int_{a}^{b}\left|{(\tau\circ\beta)^{\prime}}\right|\mathop{}\!\mathrm{d}s=\sum_{i=0}^{N-1}\int_{s_{i}}^{s_{i+1}}\left|{(\tau\circ\beta_{i})^{\prime}}\right|\mathop{}\!\mathrm{d}s.$
Proof.
item˜1 follows directly from the definition of null length and the triangle inequality.
item˜2 follows from the triangle inequality and the strict monotonicity of $\tau\circ\beta$ on each subinterval $s_{i},s_{i+1}]$, which implies that its maximum and minimum over $[a,b]$ are attained at some of the breakpoints.
item˜3 follows from
$\hat{L}_{\tau}(\beta)$ $=\sum_{i=0}^{N-1}\left|{\tau(\beta(s_{i+1}))-\tau(\beta(s_{i}))}\right|$
$=\sum_{i=0}^{N-1}\left|{\int_{s_{i}}^{s_{i+1}}(\tau\circ\beta_{i})^{\prime}\mathop{}\!\mathrm{d}s}\right|$
$=\sum_{i=0}^{N-1}\int_{s_{i}}^{s_{i+1}}\left|{(\tau\circ\beta_{i})^{\prime}}\right|\mathop{}\!\mathrm{d}s,$
where the last equality holds since $(\tau\circ\beta_{i})^{\prime}$ does not change sign on $s_{i},s_{i+1}]$, due to the causal character of $\beta_{i}$. ∎
The following proposition follows readily from the definition of null distance and Lemma 3.3.
Proposition 3.5.
Let $\tau$ be a time function on $M$. Then $\hat{d}_{\tau}$ is a semi-metric on $M$.
In general, $\hat{d}_{\tau}$ is not a metric, as it may vanish for distinct points (See [1, Prop. 3.4]). We will later introduce the local anti-Lipschitz condition on the time function, which is equivalent to the definiteness of $\hat{d}_{\tau}$. For now, we observe that the following lemma shows that non-definiteness can only occur for pairs of points lying on the same level set of $\tau$.
Lemma 3.6.
Let $\tau$ be a time function on $M$. Then, for every $p,q\in M$,
$\hat{d}_{\tau}(p,q)\geq\left|{\tau(q)-\tau(p)}\right|.$
Consequently, definiteness may fail only for pairs of points lying on the same level set of $\tau$:
$\hat{d}_{\tau}(p,q)=0\implies\tau(p)=\tau(q).$
Proof.
Let $\beta$ be a piecewise causal curve from $p$ to $q$. By Lemma 3.4, we have
$\hat{L}_{\tau}(\beta)\geq\left|{\tau(q)-\tau(p)}\right|.$
The conclusion follows by taking the infimum over all such curves. ∎
As a consequence of the previous lemma, we obtain the following causality property of the null distance.
Lemma 3.7.
Let $\tau$ be a time function on $M$. Then $\hat{d}_{\tau}$ satisfies the causality property:
$p\leq q\implies\hat{d}_{\tau}(p,q)=\tau(q)-\tau(p).$
Proof.
Suppose $p\leq q$, and let $\beta$ be a future-directed causal curve from $p$ to $q$. Since $\hat{L}_{\tau}(\beta)=\tau(q)-\tau(p)$ by Lemma 3.4, it follows that $\hat{d}_{\tau}(p,q)\leq\tau(q)-\tau(p)$. The reverse inequality follows from Lemma 3.6. ∎
3.3 Topology
Lemma 3.8.
Let $\tau$ be a time function on $M$. Then:
  1. (1)
    $\tau$ is bounded on diamonds: $p\leq x\leq q\implies\tau(p)\leq\tau(x)\leq\tau(q)$.
  2. (2)
    $\hat{d}_{\tau}$ is bounded on diamonds: $p\leq x,y\leq q\implies\hat{d}_{\tau}(x,y)\leq 2\left(\tau(q)-\tau(p)\right)$.
Proof.
item˜1 follows directly from the fact that $\tau$ is strictly increasing along future-directed causal curves.
To prove item˜2, let $\beta$ be a piecewise causal curve consisting of a future-directed causal curve from $x$ to $q$, followed by a past-directed causal curve from $q$ to $y$. Then
$\hat{L}_{\tau}(\beta)=\left|{\tau(q)-\tau(x)}\right|+\left|{\tau(y)-\tau(q)}\right|\leq 2\left(\tau(q)-\tau(p)\right),$
where the inequality follows from item˜1. Since $\hat{d}_{\tau}(x,y)$ is defined as the infimum of $\hat{L}_{\tau}$ over all piecewise causal curves from $x$ to $y$, the claim follows. ∎
Lemma 3.9.
Let $\tau$ be a time function on $M$. Then $\hat{d}_{\tau}$ is continuous on $M\times M$.
Proof.
Fix any two points $x,y\in M$. Let $\alpha_{x}$ be a future-directed timelike curve with $\alpha_{x}(0)=x$, and let $\alpha_{y}$ be a future-directed timelike curve with $\alpha_{y}(0)=y$. Choose $\delta>0$ sufficiently small so that both curves are defined on the interval $\delta,\delta]$. Then, for all $x^{\prime}\in I^{+}(\alpha_{x}(-\delta))\cap I^{-}(\alpha_{x}(\delta))$ and $y^{\prime}\in I^{+}(\alpha_{y}(-\delta))\cap I^{-}(\alpha_{y}(\delta))$, we have
$\left|{\hat{d}_{\tau}(x,y)-\hat{d}_{\tau}(x^{\prime},y^{\prime})}\right|$ $\leq\hat{d}_{\tau}(x,x^{\prime})+\hat{d}_{\tau}(y,y^{\prime})$
$\leq 2\left(\tau(\alpha_{x}(\delta))-\tau(\alpha_{x}(-\delta))\right)+2\left(\tau(\alpha_{y}(\delta))-\tau(\alpha_{y}(-\delta))\right),$
where the first inequality follows from the reverse triangle inequality, and the second from point item˜2 of Lemma 3.8. Since $\tau$ is continuous, the right-hand side tends to zero as $\delta\to 0$. Hence, $\hat{d}_{\tau}$ is continuous at $(x,y)$. ∎
As a semi-metric, $\hat{d}_{\tau}$ induces a topology on $M$ generated by the open null balls:
$\hat{B}_{r}(x)=\{y\in M\mid\hat{d}_{\tau}(x,y)<r\}.$
Proposition 3.10.
Let $\tau$ be a time function on $M$. If $\hat{d}_{\tau}$ is definite, then the topology induced by $\hat{d}_{\tau}$ coincides with the manifold topology.
Proof.
Since $\hat{d}_{\tau}$ is continuous on $M\times M$, every open null ball $\hat{B}_{r}(x)$ is open in the manifold topology. Hence, the topology induced by $\hat{d}_{\tau}$ is coarser than or equal to the manifold topology.
Conversely, suppose $U\subseteq M$ is open in the manifold topology, and fix any point $x_{0}\in U$. Choose a Riemannian metric $h$ on $M$, and denote by $d_{h}$ the corresponding distance function. Since $U$ is open, there exists $\varepsilon>0$ such that the Riemannian ball
$B=B^{h}_{\varepsilon}(x_{0})=\{y\in M\mid d_{h}(x_{0},y)<\varepsilon\}$
is contained in $U$. Since $\hat{d}_{\tau}$ is definite and continuous, and $\partial B$ is compact, the continuous function
$z\mapsto\hat{d}_{\tau}(x_{0},z)$
attains a positive minimum on $\partial B$. Define
$\varepsilon_{0}=\min_{z\in\partial B}\hat{d}_{\tau}(x_{0},z)>0.$
Now let $y\notin B$, and let $\beta$ be any piecewise causal curve from $x_{0}$ to $y$. Let $z_{0}\in\partial B$ be the first point at which $\beta$ intersects $\partial B$, and denote by $\beta_{0}$ the initial segment of $\beta$ from $x_{0}$ to $z_{0}$. Then
$\hat{L}_{\tau}(\beta)\geq\hat{L}_{\tau}(\beta_{0})\geq\varepsilon_{0}.$
Taking the infimum over all such curves $\beta$ yields
$y\notin B\implies\hat{d}_{\tau}(x_{0},y)\geq\varepsilon_{0}.$
Thus, we have
$\hat{B}_{\varepsilon_{0}}(x_{0})\subseteq B\subseteq U.$
Since $x_{0}\in U$ was arbitrary, $U$ is open in the topology induced by $\hat{d}_{\tau}$. Therefore, the two topologies coincide. ∎
3.4 Definiteness
3.4.1 Definiteness and Anti-Lipschitz Condition
Definition 3.11.
A function $f\colon M\to\mathbb{R}$ is said to be anti-Lipschitz on a subset $U\subseteq M$ if there exists a metric $d_{U}$ on $U$ such that, for all $x,y\in U$,
$x\leq y\implies f(y)-f(x)\geq d_{U}(x,y).$
The function $f$ is called locally anti-Lipschitz if it is anti-Lipschitz on some open neighborhood of every point in $M$.
Since all metrics on a manifold are locally Lipschitz equivalent, the following lemma follows immediately.
Lemma 3.12.
Let $h$ be any Riemannian metric on $M$, and denote by $d_{h}$ the corresponding distance function. Then the following conditions on a function $f\colon M\to\mathbb{R}$ are equivalent:
  1. (1)
    $f$ is locally anti-Lipschitz.
  2. (2)
    For every point $p\in M$, there exist an open neighborhood $U$ of $p$ and a constant $C>0$ such that, for all $x,y\in U$,
    $x\leq y\implies f(y)-f(x)\geq Cd_{h}(x,y).$
Proposition 3.13.
Let $\tau$ be a time function on $M$. Then $\hat{d}_{\tau}$ is definite if and only if $\tau$ is locally anti-Lipschitz.
Proof.
Suppose first that $\hat{d}_{\tau}$ is definite. Then, by Lemma 3.7, it follows that $\tau$ is anti-Lipschitz on all of $M$, and in particular locally anti-Lipschitz.
Conversely, assume that $\tau$ is locally anti-Lipschitz. Let $p,q\in M$ be distinct points, and let $U$ be an open neighborhood of $p$ on which $\tau$ is anti-Lipschitz. Choose a precompact open set $B\subseteq U$ such that $p\in B$ and $q\notin B$, and let $\beta$ be any piecewise causal curve from $p$ to $q$. Let $z_{0}\in\partial B$ be the first point at which $\beta$ intersects $\partial B$, and denote by $\beta_{0}$ the initial segment of $\beta$ from $p$ to $z_{0}$, with breakpoints $p=x_{0},x_{1},\dots,x_{N}=z_{0}$. Since $\beta_{0}\subseteq U$ and $\tau$ is anti-Lipschitz on $U$, we have
$\hat{L}_{\tau}(\beta)$ $\geq\hat{L}_{\tau}(\beta_{0})=\sum_{i=0}^{N-1}\left|{\tau(x_{i+1})-\tau(x_{i})}\right|$
$\geq\sum_{i=0}^{N-1}d_{U}(x_{i+1},x_{i})$
$\geq d_{U}(p,z_{0})$
$\geq\operatorname{dist}(p,\partial B).$
Taking the infimum over all such curves $\beta$, it follows that
$\hat{d}_{\tau}(p,q)\geq\operatorname{dist}(p,\partial B)>0,$
which proves that $\hat{d}_{\tau}$ is definite. ∎
Combining Propositions 3.53.10, and 3.13, we obtain the following result.
Theorem 3.14.
Let $\tau$ be a time function on $M$. If $\tau$ is locally anti-Lipschitz, then $\hat{d}_{\tau}$ is a metric on $M$ that induces the manifold topology.
3.4.2 Temporal Functions
Every temporal function $f\colon M\to\mathbb{R}$ is, in particular, a time function. As such, the associated null distance $\hat{d}_{f}$ is a semi-metric on $M$. A natural question is whether $\hat{d}_{f}$ is actually a metric, or equivalently, whether $f$ is locally anti-Lipschitz. This is indeed the case for $C^{1}$ temporal functions. Moreover, within the $C^{1}$ class, the local anti-Lipschitz condition characterizes temporal functions.
Theorem 3.15.
Let $f\colon M\to\mathbb{R}$ be a $C^{1}$ function. Then $f$ is locally anti-Lipschitz if and only if it is a temporal function.
The following proposition provides an equivalent characterization of the local anti-Lipschitz condition for $C^{1}$ functions.
Proposition 3.16.
Let $h$ be any Riemannian metric on $M$, and denote by $d_{h}$ the corresponding distance function. Then the following conditions on a $C^{1}$ function $f\colon M\to\mathbb{R}$ are equivalent:
  1. (1)
    For every point $p\in M$, there exist an open neighborhood $U$ of $p$ and a constant $C>0$ such that, for all $x,y\in U$,
    $x\leq y\implies f(y)-f(x)\geq Cd_{h}(x,y).$
  2. (2)
    For every point $p\in M$, there exist an open neighborhood $U$ of $p$ and a constant $C>0$ such that, for all future-directed causal vectors $X\in TU$,
    $g(\nabla f,X)\geq C\lVert{X}\rVert_{h}.$
Proof.
[…]
$f(\alpha(t))-f(\alpha(0))\geq Ct.$
Dividing both sides by $t$ and taking the limit $t\to 0^{+}$ yields
$(f\circ\alpha)^{\prime}(0)=\lim_{t\to 0^{+}}\frac{f(\alpha(t))-f(\alpha(0))}{t}\geq C.$
Since $(f\circ\alpha)^{\prime}(0)=g(\nabla f,X/\lVert{X}\rVert_{h})$, we conclude that
$g(\nabla f,X)\geq C\lVert{X}\rVert_{h}.$
By continuity of $g$ and $h$, this inequality extends to all future-directed causal vectors $X\in TU$.
[…]
$f(y)-f(x)=\int_{a}^{b}(f\circ\alpha)^{\prime}\mathop{}\!\mathrm{d}s=\int_{a}^{b}g(\nabla f,\alpha^{\prime})\mathop{}\!\mathrm{d}s\geq CL_{h}(\alpha)\geq Cd_{h}(x,y),$
where the first inequality follows from the assumption in item˜2, and the second from the definition of $d_{h}(x,y)$ as the infimum of the $h$-length over all smooth curves from $x$ to $y$. ∎
Definition 3.17.
Let $h$ be any Riemannian metric on $M$ and let $\mathcal{T}\subseteq TM$ be a collection of tangent vectors. We say that $\mathcal{T}$ is locally bounded away from zero if, for every point $p\in M$, there exist an open neighborhood $U$ of $p$ and a constant $C>0$ such that, for all $T\in\mathcal{T}\cap TU$, $\lVert{T}\rVert_{g}\geq C$.
Lemma 3.18.
Let $h$ be any Riemannian metric on $M$, and let $\mathcal{T}\subseteq TM$ be a collection of tangent vectors. If the vectors in $\mathcal{T}$ are past-directed timelike and locally bounded away from zero, then, for every point $p\in M$, there exist an open neighborhood $U$ of $p$ and a constant $C>0$ such that, for all future-directed causal vectors $X\in TU$ and all $T\in\mathcal{T}\cap TU$,
$g(T,X)\geq C\lVert{X}\rVert_{h}.$
Proof.
Suppose, for contradiction, that the conclusion fails at some point $p_{0}\in M$. Let $U$ and $C$ be as in Definition 3.17, corresponding to $p=p_{0}$. Then there exist a sequence of points $p_{j}\in U$, with $p_{j}\to p_{0}$, a sequence of future-directed causal vectors $X_{j}\in T_{p_{j}}M$, and a sequence of past-directed timelike vectors $T_{j}\in\mathcal{T}\cap T_{p_{j}}M$ such that
$g(T_{j},X_{j})<j^{-1}\lVert{X_{j}}\rVert_{h}.$
By rescaling and passing to a subsequence if necessary, we may assume that $\lVert{X_{j}}\rVert_{h}=1$ for all $j$ and that $X_{j}\to X_{0}$ for some future-directed causal vector $X_{0}\in T_{p_{0}}M$. Assume first that the sequence $T_{j}$ is bounded in the $h$-norm. Then, up to a further subsequence, $T_{j}\to V_{0}\in T_{p_{0}}M$, where $V_{0}$ is either past-directed causal or zero. Passing to the limit in the inequality yields
$g(V_{0},X_{0})=0.$
By Lemma 2.3, it follows that $V_{0}$ is either null or zero. In both cases, we have $\lVert{T_{j}}\rVert_{g}\to\lVert{V_{0}}\rVert_{g}=0$, contradicting the assumption $\lVert{T_{j}}\rVert_{g}\geq C>0$. Suppose instead that $\lVert{T_{j}}\rVert_{h}\to\infty$. Define $W_{j}=T_{j}/\lVert{T_{j}}\rVert_{h}$. Then, for sufficiently large $j$, we have $\lVert{T_{j}}\rVert_{h}\geq 1$, and thus
$g(W_{j},X_{j})<j^{-1}\lVert{T_{j}}\rVert_{h}^{-1}\leq j^{-1}.$
Since $\lVert{W_{j}}\rVert_{h}=1$ for all $j$, we may extract a subsequence such that $W_{j}\to W_{0}$ for some past-directed causal vector $W_{0}\in T_{p_{0}}M$. Passing to the limit in the inequality yields
$g(W_{0},X_{0})=0.$
By Lemma 2.3, it follows that $V_{0}$ is null. This contradicts the assumption, since for sufficiently large $j$ we have
$0<C\leq\frac{\lVert{T_{j}}\rVert_{g}}{\lVert{T_{j}}\rVert_{h}}=\lVert{W_{j}}\rVert_{g}\to\lVert{W_{0}}\rVert_{g}=0.$
We are now ready to prove Theorem 3.15.
Proof of Theorem 3.15.
Suppose first that $f$ is locally anti-Lipschitz, and let $h$ be any Riemannian metric on $M$. Then, by Proposition 3.16 and Lemma 3.12, for every point $p\in M$, there exist an open neighborhood $U$ of $p$ and a constant $C>0$ such that, for all future-directed causal vectors $X\in TU$,
$g(\nabla f,X)\geq C\lVert{X}\rVert_{h}.$
By Lemma 2.3, it follows that $\nabla f$ is past-directed timelike on all of $M$, and therefore $f$ is a temporal function.
Conversely, suppose that $f$ is a temporal function. Then $\nabla f$ is a continuous past-directed timelike vector field on $M$. By continuity, the collection $\{(\nabla f)_{p}\}_{p\in M}$ is locally bounded away from zero. It then follows from Lemma 3.18, Proposition 3.16, and Lemma 3.12 that $f$ is locally anti-Lipschitz. ∎
4 Main Result
The purpose of this section is to establish Theorem 1.2. We begin by analyzing a guiding example in the two-dimensional Minkowski spacetime.
Example 4.1.
Consider the temporal function $f(t,x)=t$ on the two-dimensional Minkowski spacetime $\mathbb{M}^{2}$. For any fixed $t\in\mathbb{R}$, let $p,q\in M_{t}$ be two distinct points lying on the same level set
$M_{t}=f^{-1}(t)=\{(t^{\prime},x^{\prime})\in\mathbb{R}^{2}\mid t^{\prime}=t\},$
and let $d=\left|{x_{q}-x_{p}}\right|$ denote the Euclidean distance between their spatial components. Since the induced Riemannian metric $h_{t}$ on $M_{t}$ coincides with the standard Euclidean metric, it follows that
$d_{h_{t}}(p,q)=d.$
Now consider the point
$r=\left(t-\frac{d}{2},\frac{x_{p}+x_{q}}{2}\right),$
which lies at the intersection of the past null cones of $p$ and $q$. Define the piecewise causal curve $\beta$ as the concatenation of the null segments $[p,r]$ and $[r,q]$. Then
$\hat{L}_{f}(\beta)=2\left|{t_{r}-t}\right|=d.$
Therefore,
$\frac{\hat{d}_{f}(p,q)}{d_{h_{t}}(p,q)}\leq\frac{\hat{L}_{f}(\beta)}{d_{h_{t}}(p,q)}=1.$
The key idea in the previous example is the construction of a piecewise causal curve whose null length equals the spatial distance between its endpoints. The strategy underlying the proof of Theorem 1.2 is to asymptotically reproduce this construction, as formalized in the following lemma.
Lemma 4.2.
Let $(M,g)$ be a spacetime admitting a smooth temporal function $f\colon M\to\mathbb{R}$. Suppose there exist $t\in\mathbb{R}$ and $C>0$ such that the level set $M_{t}=f^{-1}(t)$ is nonempty and the equality
$g(\nabla f,\nabla f)=-C^{2}$
holds at every point of $M_{t}$. Fix a point $p\in M_{t}$, and let $\gamma\colon[0,L]\to M_{t}$ be a smooth unit-speed curve starting at $p$. Then, for every $\varepsilon\in(0,1)$ and every $s\in(0,L]$, there exists a piecewise causal curve $\beta_{s}$ from $p$ to $\gamma(s)$ such that [^3] The curve $\beta_{s}$ depends on both $s$ and $\varepsilon$, but only the dependence on $s$ is made explicit to simplify the notation.
$\lim_{\varepsilon\to 0^{+}}\lim_{s\to 0^{+}}\frac{\hat{L}_{f}(\beta_{s})}{L_{h_{t}}({\gamma}|_{[0,s]})}=C,$
where $h_{t}$ denotes the induced Riemannian metric on $M_{t}$ and $L_{h_{t}}$ is the associated length functional. Consequently,
$\lim_{s\to 0^{+}}\frac{\hat{d}_{f}(p,\gamma(s))}{L_{h_{t}}({\gamma}|_{[0,s]})}\leq C.$
To prove this lemma, we first establish a preliminary result showing that, within a sufficiently small normal coordinate neighborhood, piecewise causal curves for a suitably perturbed Minkowski metric are mapped to piecewise causal curves for the spacetime metric.
For each $\varepsilon\in(0,1)$, define the $\varepsilon$-slim Minkowski metric on $\mathbb{R}^{n+1}$ by
$\eta_{\varepsilon}=-(1-\varepsilon)dt^{2}+d(x^{1})^{2}+\dots+d(x^{n})^{2},$
and denote by $C_{\varepsilon}$ the corresponding causal cone:
$C_{\varepsilon}=\{v\in\mathbb{R}^{n+1}\setminus\{0\}\mid\eta_{\varepsilon}(v,v)\leq 0\}.$
We then have the following result.
Lemma 4.3.
Let $(U,\varphi)$ be a normal coordinate chart centered at $p\in M$. Then, for every $\varepsilon\in(0,1)$, there exists an open neighborhood $U_{\varepsilon}\subseteq U$ of $p$ such that every piecewise $\eta_{\varepsilon}$-causal curve in $\varphi(U_{\varepsilon})$ is mapped by $\varphi^{-1}$ to a piecewise $g$-causal curve in $U_{\varepsilon}$.
Proof.
At the origin $0\in\mathbb{R}^{n+1}$, the pullback of $g$ under $\varphi^{-1}$ satisfies [^4] We identify the tangent space at any point of $\mathbb{R}^{n+1}$ with $\mathbb{R}^{n+1}$ itself.
$((\varphi^{-1})^{*}g)_{0}(v,v)=\eta(v,v)<\eta_{\varepsilon}(v,v)$
for all $v\in C_{\varepsilon}$. In particular, the inequality holds on the compact set
$D_{\varepsilon}=S^{n}\cap C_{\varepsilon},$
where $S^{n}$ denotes the unit sphere in $\mathbb{R}^{n+1}$. Define the function
$F\colon\varphi(U)\times D_{\varepsilon}\to\mathbb{R},\quad F(x,v)=((\varphi^{-1})^{*}g)_{x}(v,v)-\eta_{\varepsilon}(v,v).$
By continuity of $F$ and the fact that $F(0,v)<0$ for all $v\in D_{\varepsilon}$, for every $v\in D_{\varepsilon}$ there exist open neighborhoods $U_{v}\subseteq\varphi(U)$ of 0 and $V_{v}\subseteq D_{\varepsilon}$ of $v$ such that $F(x,w)<0$ for all $(x,w)\in U_{v}\times V_{v}$. By compactness of $D_{\varepsilon}$, there exist finitely many points $v_{1},\dots,v_{N}\in D_{\varepsilon}$ such that
$D_{\varepsilon}\subseteq\bigcup_{i=1}^{N}V_{v_{i}}.$
For each $i=1,\dots,N$, let $r_{i}>0$ be such that the Euclidean open ball $B_{r_{i}}(0)$ is contained in $U_{v_{i}}$, and define
$r_{\min}=\min_{1\leq i\leq N}r_{i},\quad U_{\varepsilon}=\varphi^{-1}(B_{r_{\min}}(0)).$
Fix $x\in\varphi(U_{\varepsilon})$ and $w\in C_{\varepsilon}$. Then the normalized vector $\hat{w}=w/\left|{w}\right|$ lies in $D_{\varepsilon}$, hence in some $V_{v_{i}}$, and $x\in B_{r_{i}}(0)\subseteq U_{v_{i}}$. Therefore,
$F(x,w)=\left|{w}\right|^{2}F(x,\hat{w})<0,$
which implies
$((\varphi^{-1})^{*}g)_{x}(w,w)<\eta_{\varepsilon}(w,w).$
Thus, $\mathrm{d}_{x}(\varphi^{-1})$ maps $\eta_{\varepsilon}$-causal vectors to $g$-causal vectors. Consequently, $\varphi^{-1}$ maps piecewise $\eta_{\varepsilon}$-causal curves in $\varphi(U_{\varepsilon})$ to piecewise $g$-causal curves in $U_{\varepsilon}$. ∎
We now proceed with the proof of Lemma 4.2.
Proof of Lemma 4.2.
Let $(U,\varphi)$ be a normal coordinate chart centered at $p$. Since $(\nabla f)_{p}$ is a past-directed timelike vector satisfying
$g_{p}((\nabla f)_{p},(\nabla f)_{p})=-C^{2},$
and $\gamma^{\prime}(0)$ is a spacelike unit vector, we may, up to a Lorentz transformation, assume that
$\mathrm{d}_{p}\varphi((\nabla f)_{p})=-Ce_{0}$ (4.4)
and
$\mathrm{d}_{p}\varphi(\gamma^{\prime}(0))=e_{1},$ (4.5)
where $\{e_{i}\}_{i=0}^{n}$ denotes the standard orthonormal basis of the Minkowski spacetime $\mathbb{M}^{n+1}$. Fix $\varepsilon\in(0,1)$, and let $U_{\varepsilon}\subseteq U$ be the open neighborhood provided by Lemma 4.3. For each $s\in(0,L]$, define $q_{s}=\varphi(\gamma(s))$ and write $q_{s}=(t_{s},x_{s})$, where $t_{s}\in\mathbb{R}$ and $x_{s}\in\mathbb{R}^{n}$ denote the temporal and spatial components of $q_{s}$, respectively. Equation (4.5) ensures that $\left|{x_{s}}\right|\neq 0$ for sufficiently small $s$. Furthermore, the first-order Taylor expansion of $\varphi\circ\gamma$ at zero yields
$\left|{x_{s}}\right|=s+o(s),\quad t_{s}=o(s)\quad\text{as }\,s\to 0^{+}.$ (4.6)
It follows that, for sufficiently small $s$,
$\left|{x_{s}}\right|^{2}-(1-\varepsilon)t_{s}^{2}\neq 0,$
ensuring that the quantity
$t^{*}_{s}=\frac{\left|{x_{s}}\right|^{2}-(1-\varepsilon)t_{s}^{2}}{2\sqrt{1-\varepsilon}\left(\left|{x_{s}}\right|-\sqrt{1-\varepsilon}\,t_{s}\right)}.$
is well-defined and nonzero. Consequently, for sufficiently small $s$, the point
$r_{s}=\left(t^{*}_{s},\sqrt{1-\varepsilon}\,t^{*}_{s}\frac{x_{s}}{\left|{x_{s}}\right|}\right),$
is well-defined and nonzero. Substituting the expansions in (4.6) into the expression for $t^{*}_{s}$ and simplifying by absorbing higher-order terms into $o(s)$, we find
$t^{*}_{s}=\frac{1}{2\sqrt{1-\varepsilon}}(s+o(s))\quad\text{as }\,s\to 0^{+},$
and thus,
$\lim_{s\to 0^{+}}\frac{t^{*}_{s}}{s}=\frac{1}{2\sqrt{1-\varepsilon}}.$
A direct computation shows that the straight-line segments $0,r_{s}]$ and $r_{s},q_{s}]$ are $\eta_{\varepsilon}$-null. Furthermore, the Euclidean norm of $r_{s}$ satisfies
$\left|{r_{s}}\right|=\sqrt{2-\varepsilon}\,\left|{t^{*}_{s}}\right|=\frac{1}{2}\sqrt{\frac{2-\varepsilon}{1-\varepsilon}}\,\left|{s+o(s)}\right|\quad\text{as }\,s\to 0^{+}.$ (4.7)
Define the piecewise $\eta_{\varepsilon}$-null curve $\tilde{\beta}_{s}$ as the concatenation of the segments $0,r_{s}]$ and $r_{s},q_{s}]$. Since $\varphi(U_{\varepsilon})$ is open and both $r_{s}$ and $q_{s}$ converge to 0 as $s\to 0^{+}$, the curve $\tilde{\beta}_{s}$ is entirely contained in $\varphi(U_{\varepsilon})$ for sufficiently small $s$. By Lemma 4.3, the curve $\beta_{s}=\varphi^{-1}\circ\tilde{\beta}_{s}$ is piecewise $g$-causal in $U_{\varepsilon}$, and its null length satisfies
$\hat{L}_{f}(\beta_{s})$ $=2\left|{(f\circ\varphi^{-1})(r_{s})-(f\circ\varphi^{-1})(0)}\right|$
$=2\left|{(\nabla(f\circ\varphi^{-1}))_{0}\cdot r_{s}+o(\left|{r_{s}}\right|)}\right|\quad\text{as }\,\left|{r_{s}}\right|\to 0,$
where the second equality follows from the first-order Taylor expansion of $f\circ\varphi^{-1}$ at the origin. Moreover, using the definition of the gradient, equation (4.4), and the fact that $\varphi$ is a normal coordinate chart, we obtain
$(\nabla(f\circ\varphi^{-1}))_{0}\cdot r_{s}$ $=\mathrm{d}_{0}(f\circ\varphi^{-1})(r_{s})=\mathrm{d}_{p}f(\mathrm{d}_{0}\varphi^{-1}(r_{s}))$
$=g_{p}((\nabla f)_{p},\mathrm{d}_{0}\varphi^{-1}(r_{s}))=\eta(\mathrm{d}_{p}\varphi((\nabla f)_{p}),r_{s})$
$=\eta(-Ce_{0},r_{s})=Ct^{*}_{s}.$
It follows that
$\hat{L}_{f}(\beta_{s})=2\left|{Ct^{*}_{s}+o(\left|{r_{s}}\right|)}\right|\quad\text{as }\,\left|{r_{s}}\right|\to 0.$
Since $\gamma$ is unit-speed, we have $L_{h_{t}}({\gamma}|_{[0,s]})=s$. Therefore,
$\frac{\hat{L}_{f}(\beta_{s})}{L_{h_{t}}({\gamma}|_{[0,s]})}=\frac{2\left|{Ct^{*}_{s}+o(\left|{r_{s}}\right|)}\right|}{s}\quad\text{as }\,\left|{r_{s}}\right|\to 0.$ (4.8)
Using (4.7), we deduce that
$\lim_{s\to 0^{+}}\frac{o(\left|{r_{s}}\right|)}{s}=0.$
Taking the limits $s\to 0^{+}$ and then $\varepsilon\to 0^{+}$ in (4.8), we conclude that
$\lim_{\varepsilon\to 0^{+}}\lim_{s\to 0^{+}}\frac{\hat{L}_{f}(\beta_{s})}{L_{h_{t}}({\gamma}|_{[0,s]})}=\lim_{\varepsilon\to 0^{+}}\lim_{s\to 0^{+}}\frac{2\left|{Ct^{*}_{s}+o(\left|{r_{s}}\right|)}\right|}{s}=\lim_{\varepsilon\to 0^{+}}\frac{C}{\sqrt{1-\varepsilon}}=C.$
We now complete the argument by establishing Theorem 1.2.
Proof of Theorem 1.2.
Let $\delta>0$, and let $\gamma:[0,L]\to M_{t}$ be a smooth unit-speed curve from $p$ to $q$ satisfying
$L_{h_{t}}(\gamma)<d_{h_{t}}(p,q)+\delta.$ (4.9)
By Lemma 4.2, for every $s\in[0,L]$, there exists $\varepsilon_{s}>0$ such that, for all $s^{\prime}\in[0,L]$ with $\left|{s-s^{\prime}}\right|<\varepsilon_{s}$, we have
$\hat{d}_{f}(\gamma(s),\gamma(s^{\prime}))\leq(C+\delta)L_{h_{t}}({\gamma}|_{[s,s^{\prime}]}).$
The collection of intervals $\{(s-\varepsilon_{s}/2,s+\varepsilon_{s}/2)\}_{s\in[0,L]}$ forms an open cover of $[0,L]$. Using the Lebesgue’s number lemma, we can find a partition $0=s_{0}<\dots<s_{N}=L$ such that, for each $i=0,\dots,N-1$,
$\hat{d}_{f}(\gamma(s_{i}),\gamma(s_{i+1}))\leq(C+\delta)L_{h_{t}}({\gamma}|_{[s_{i},s_{i+1}]}).$ (4.10)
Applying the triangle inequality together with estimates (4.10) and (4.9), we obtain
$\hat{d}_{f}(p,q)$ $\leq\sum_{i=0}^{N-1}\hat{d}_{f}(\gamma(s_{i}),\gamma(s_{i+1}))$
$\leq(C+\delta)\sum_{i=0}^{N-1}L_{h_{t}}({\gamma}|_{[s_{i},s_{i+1}]})$
$=(C+\delta)L_{h_{t}}(\gamma)$
$<(C+\delta)(d_{h_{t}}(p,q)+\delta).$
The conclusion follows by letting $\delta\to 0$. ∎
As a direct consequence of Theorem 1.2 and Example 2.9, we obtain the following result.
Corollary 4.11.
In a GRW spacetime
$(M,g)=(I\times S,-dt^{2}+f(t)^{2}h),$
consider a smooth temporal function of the form $\tau(t,x)=\phi(t)$, where $\phi\colon I\to\mathbb{R}$ satisfies $\phi^{\prime}(t)>0$ for all $t\in I$. For any $t\in I$, consider the level set $M_{t}=\tau^{-1}(t)$. Then, for every $p,q\in M_{t}$,
$\hat{d}_{\tau}(p,q)\leq\phi^{\prime}(t)d_{h_{t}}(p,q),$
where $h_{t}$ denotes the induced Riemannian metric on $M_{t}$ and $d_{h_{t}}$ is the corresponding distance function.
5 Cosmological Time
This section focuses on the notion of regular cosmological time, as introduced by Andersson, Galloway, and Howard in [2]. It also presents the proof of Theorem 1.3.
5.1 Regular Cosmological Time
The cosmological time function $\tau_{g}\colon M\to[0,\infty]$ is defined by
$\tau_{g}(q)=\sup_{p\leq q}d_{g}(p,q).$
In general, the function $\tau_{g}$ is not necessarily well-behaved; for instance, it is identically infinite in Minkowski spacetime. In [2], Andersson, Galloway, and Howard introduced a notion of regularity for the cosmological time function and proved that, under this condition, several important properties hold. A future-directed causal curve $\gamma\colon(a,b)\to M$ is said to be past inextendible if there exists no point $p\in M$ such that
$\lim_{s\to a^{+}}\gamma(s)=p.$
The cosmological time function $\tau_{g}$ is called regular if it satisfies the following conditions:
  1. (1)
    $\tau_{g}(q)<\infty$ for all $q\in M$;
  2. (2)
    $\tau_{g}\to 0$ along all past inextendible future-directed causal curves.
Proposition 5.1 ([2, Thm. 1.2]).
Suppose that the cosmological time function $\tau_{g}$ is regular. Then the following properties hold:
  1. (1)
    $\tau_{g}$ is a time function, and for all $x,y\in M$,
    $x\leq y\implies\tau_{g}(y)-\tau_{g}(x)\geq d_{g}(x,y).$
  2. (2)
    The gradient $\nabla\tau_{g}$ exists almost everywhere.
  3. (3)
    For every point $q\in M$, there exists a future-directed timelike unit-speed geodesic $\gamma_{q}\colon(0,\tau_{g}(q)]\to M$ such that $\gamma_{q}(\tau_{g}(q))=q$ and $\tau_{g}(\gamma_{q}(t))=t$ for all $t\in(0,\tau_{g}(q))$.
A curve as in point item˜3 above is called a generator at $q$. The following result shows that, wherever it exists, the gradient of a regular $\tau_{g}$ is a past-directed timelike unit vector.
Proposition 5.2.
Suppose that the cosmological time function $\tau_{g}$ is regular and differentiable at a point $q\in M$. Then there exists a unique generator $\gamma_{q}$ at $q$, and $(\nabla\tau_{g})_{q}=-\gamma_{q}^{\prime}(\tau_{g}(q))$.
Proof.
Set $\tau=\tau_{g}$, and let $\gamma=\gamma_{q}$ be any generator at $q$. Define $u=\gamma^{\prime}(\tau(q))$, which is a future-directed timelike unit vector at $q$. Then the gradient of $\tau$ at $q$ can be expressed as
$(\nabla\tau)_{q}=-au+bv,$
where $v\in T_{q}M$ is a spacelike unit vector orthogonal to $u$, and $a,b\geq 0$. Since $\gamma$ is a generator at $q$, we have
$a=g((\nabla\tau)_{q},u)=\frac{\mathrm{d}}{\mathrm{d}t}(\tau\circ\gamma)=\frac{\mathrm{d}}{\mathrm{d}t}(t)=1.$
Thus, $(\nabla\tau)_{q}=-u+bv$. By Lemma 2.7, the vector $(\nabla\tau)_{q}$ must be causal, which implies $0\leq b\leq 1$. In particular,
$\lVert{(\nabla\tau)_{q}}\rVert_{g}=\sqrt{1-b^{2}}\leq 1.$
Now fix any future-directed timelike unit vector $V\in T_{q}M$, and let $\alpha\colon(-\delta,\delta)\to M$ be a future-directed timelike unit-speed curve with $\alpha(0)=q$ and $\alpha^{\prime}(0)=V$. By point item˜1 of Proposition 5.1, we obtain
$g((\nabla\tau)_{q},V)=\lim_{t\to 0^{+}}\frac{\tau(\alpha(t))-\tau(\alpha(0))}{t}\geq\lim_{t\to 0^{+}}\frac{d_{g}(\alpha(0),\alpha(t))}{t}\geq 1.$
We now show that this forces $b=0$, hence $(\nabla\tau)_{q}=-u$. First, suppose $b\neq 1$, so that $(\nabla\tau)_{q}$ is not null. Applying the inequality above to $V=-(\nabla\tau)_{q}/\lVert{(\nabla\tau)_{q}}\rVert_{g}$ yields
$\lVert{(\nabla\tau)_{q}}\rVert_{g}=g((\nabla\tau)_{q},V)\geq 1.$
Combining this with the earlier bound $\lVert{(\nabla\tau)_{q}}\rVert_{g}\leq 1$, we conclude that $\lVert{(\nabla\tau)_{q}}\rVert_{g}=1$, and thus $b=0$. Now suppose $b=1$. For every $A\geq 1$, consider the future-directed timelike unit vector
$V=Au-\sqrt{A^{2}-1}\,v.$
Then
$1\leq g((\nabla\tau)_{q},V)=A-\sqrt{A^{2}-1}.$
However, the right-hand side tends to 0 as $A\to\infty$, yielding a contradiction. Therefore, the case $b=1$ is excluded, and from the previous argument we deduce that $b=0$.
Finally, the uniqueness of the generator at $q$ follows from the uniqueness of geodesics with given initial data. ∎
5.2 Initial Level Set
As a consequence of Proposition 5.2, a smooth regular cosmological time function $\tau_{g}$ is a smooth temporal function with unit gradient everywhere. Then the associated null distance $\hat{d}_{\tau_{g}}$ is a metric on $M$ that induces the manifold topology. Furthermore, by Lemma 3.6, $\tau_{g}$ is $1$-Lipschitz with respect to $\hat{d}_{\tau_{g}}$, and by Corollary 2.18, it admits a unique $1$-Lipschitz extension $\overline{\tau_{g}}\colon\overline{M}\to[0,\infty)$ to the metric completion of $(M,\hat{d}_{\tau_{g}})$.
Lemma 5.3.
Suppose that the cosmological time function $\tau_{g}$ is smooth and regular. Then, for every $q\in M$, the unique generator $\gamma_{q}\colon(0,\tau_{g}(q)]\to M$ at $q$ is $1$-Lipschitz with respect to $\hat{d}_{\tau_{g}}$. In particular, $\gamma_{q}$ admits a unique $1$-Lipschitz extension $\overline{\gamma}_{q}\colon[0,\tau_{g}(q)]\to\overline{M}$.
Proof.
For all $t,t^{\prime}\in(0,\tau_{g}(q)]$, we have
$\hat{d}_{\tau_{g}}(\gamma_{q}(t),\gamma_{q}(t^{\prime}))\leq\hat{L}_{\tau_{g}}({\gamma_{q}}|_{[t,t^{\prime}]})=\left|{\tau_{g}(t)-\tau_{g}(t^{\prime})}\right|=\left|{t-t^{\prime}}\right|$
where we used Lemma 3.4 and the fact that $\gamma_{q}$ is a generator. This shows that $\gamma_{q}$ is $1$-Lipschitz with respect to $\hat{d}_{\tau_{g}}$. The existence and uniqueness of the extension then follow from Corollary 2.18. ∎
We are now ready to prove Theorem 1.3.
Proof of Theorem 1.3.
We first prove that $\overline{\tau_{g}}^{-1}(0)\neq\emptyset$. Fix a point $q\in M$, and let $\gamma_{q}\colon(0,\tau_{g}(q)]\to M$ be the unique generator at $q$. Consider a sequence $t_{k}\in(0,\tau(q)]$ with $t_{k}\to 0$. Then the sequence $\gamma_{q}(t_{k})\in M$ is Cauchy with respect to $\hat{d}_{\tau_{g}}$:
$\hat{d}_{\tau_{g}}(\gamma_{q}(t_{k}),\gamma_{q}(t_{j}))\leq\left|{t_{k}-t_{j}}\right|\to 0,$
where we used Lemma 5.3. Since $\overline{M}$ is complete, the sequence converges to a point $q_{0}\in\overline{M}$. By continuity of $\overline{\tau_{g}}$ and the fact that $\gamma_{q}$ is a generator, we obtain
$\overline{\tau_{g}}(q_{0})=\lim_{k}\tau_{g}(\gamma_{q}(t_{k}))=\lim_{k}t_{k}=0,$
which shows that $q_{0}\in\overline{\tau_{g}}^{-1}(0)$.
We now prove that $\overline{\tau_{g}}^{-1}(0)$ consists of a single point. Since $\tau_{g}$ is a smooth temporal function with unit gradient everywhere, Theorem 1.2 implies that for any $t>0$ such that $M_{t}$ is nonempty, and for every $p,q\in M_{t}$,
$\hat{d}_{\tau_{g}}(p,q)\leq d_{h_{t}}(p,q).$
In particular, it follows that
$\lim_{t\to 0^{+}}\operatorname{diam}_{\hat{d}_{\tau_{g}}}(M_{t})=0.$
Since, for sufficiently small $t>0$, $M_{t}$ is a nonempty closed and bounded subset of $\overline{M}$, and $\overline{\tau_{g}}^{-1}(0)$ is nonempty and closed, Lemma 2.15 applies; hence, it suffices to show that $M_{t}$ converges to $\overline{\tau_{g}}^{-1}(0)$ in the Hausdorff sense as $t\to 0$. Indeed, in that case,
$\operatorname{diam}(\overline{\tau_{g}}^{-1}(0))=\lim_{t\to 0^{+}}\operatorname{diam}_{\hat{d}_{\tau_{g}}}(M_{t})=0,$
which implies that $\overline{\tau_{g}}^{-1}(0)$ consists of a single point. For every $t>0$ such that $M_{t}$ is nonempty, we have
$d_{H}(M_{t},\overline{\tau_{g}}^{-1}(0))=\max\{\sup\nolimits_{q_{t}\in M_{t}}\operatorname{dist}(q_{t},\overline{\tau_{g}}^{-1}(0)),\ \sup\nolimits_{q_{0}\in\overline{\tau_{g}}^{-1}(0)}\operatorname{dist}(q_{0},M_{t})\}.$
We estimate each term separately. Fix any point $q_{t}\in M_{t}$, and let $\gamma_{q_{t}}\colon(0,t]\to M$ be the unique generator at $q_{t}$. Let $\overline{\gamma}_{q_{t}}\colon[0,t]\to\overline{M}$ be its $1$-Lipschitz extension. Then
$\operatorname{dist}(q_{t},\overline{\tau_{g}}^{-1}(0))\leq\hat{d}_{\tau_{g}}(q_{t},\overline{\gamma}_{q_{t}}(0))=\hat{d}_{\tau_{g}}(\overline{\gamma}_{q_{t}}(t),\overline{\gamma}_{q_{t}}(0))\leq t.$ (5.4)
Now fix any point $q_{0}\in\overline{\tau_{g}}^{-1}(0)$, and let $q_{i}\in M$ be a sequence converging to $q_{0}$. Choose any point $q\in M_{t}$, and let $\gamma_{q}\colon(0,t]\to M$ be its generator. Given $\varepsilon>0$, choose $0<\tilde{t}\leq t$ such that
$\operatorname{diam}_{\hat{d}_{\tau_{g}}}(M_{t^{\prime}})<\varepsilon/2\quad\text{for all }0<t^{\prime}\leq\tilde{t}$.
Since $q_{i}\to q_{0}$ and $\tau_{g}(q_{i})\to 0$, we can select $i$ sufficiently large so that
$\hat{d}_{\tau_{g}}(q_{0},q_{i})<\varepsilon/2\quad\text{ and }\quad\tau_{g}(q_{i})<\tilde{t}$.
Then
$\operatorname{dist}(q_{0},M_{t})$ $\leq\hat{d}_{\tau_{g}}(q_{0},q)$
$\leq\hat{d}_{\tau_{g}}(q_{0},q_{i})+\hat{d}_{\tau_{g}}(q_{i},\gamma_{q}(\tau_{g}(q_{i})))+\hat{d}_{\tau_{g}}(\gamma_{q}(\tau_{g}(q_{i})),q)$
$<\varepsilon/2+\operatorname{diam}_{\hat{d}_{\tau_{g}}}(M_{\tau_{g}(q_{i})})+(t-\tau_{g}(q_{i}))$
$<t+\varepsilon.$
Taking the limit $\varepsilon\to 0$, we conclude that
$\operatorname{dist}(q_{0},M_{t})\leq t.$ (5.5)
Combining estimates (5.4) and (5.5), we obtain
$d_{H}(M_{t},\overline{\tau_{g}}^{-1}(0))\leq t,$
which implies that $M_{t}$ converges to $\overline{\tau_{g}}^{-1}(0)$ in the Hausdorff sense as $t\to 0$. ∎
6 Appendix
6.1 Deferred Proofs
Lemma (Lemma 2.2 (restated)).
Let $X,T\in TM$ be tangent vectors at the same point, with $X$ causal and $T$ timelike. Then $g(X,T)\neq 0$.
Proof.
Suppose, for contradiction, that $g(X,T)=0$. Since $T$ is timelike and the metric $g$ has signature $(1,n)$, $X$ cannot be timelike and must therefore be null. Given that $g(X,T)=0$ and $X$ is null, equality is attained in the reverse Cauchy–Schwarz inequality, which implies that $X$ and $T$ are linearly dependent. This is a contradiction, as $X$ is null and $T$ is timelike. ∎
Lemma (Lemma 2.3 (restated)).
Let $p\in M$, and let $h_{p}$ be any Euclidean inner product on $T_{p}M$. Then the following conditions on a tangent vector $T\in T_{p}M$ are equivalent:
  1. (1)
    $T$ is past-directed timelike.
  2. (2)
    For every future-directed causal vector $X\in T_{p}M$, $g(X,T)>0$.
  3. (3)
    There exists a constant $C>0$ such that, for all future-directed causal vectors $X\in T_{p}M$,
    $g(T,X)\geq C\lVert{X}\rVert_{h_{p}}.$
Proof.
[…]
$T^{\prime}=T+\lambda\vartheta,$
where $\lambda<0$ is chosen so that
$g(X,T^{\prime})=g(X,T)+\lambda g(X,\vartheta)=0.$
Then $T^{\prime}$ is timelike and $g(X,T^{\prime})=0$, contradicting Lemma 2.2.
[…]
$g(T,X_{j})<j^{-1}\lVert{X_{j}}\rVert_{h_{p}}.$
By rescaling and passing to a subsequence if necessary, we may assume that $\lVert{X_{j}}\rVert_{h_{p}}=1$ for all $j$ and that $X_{j}\to X_{0}$ for some future-directed causal vector $X_{0}\in T_{p}M$. Passing to the limit in the inequality yields
$g(T,X_{0})=0,$
which contradicts item˜2.
[…]
$g(T,\vartheta)\geq C\lVert{\vartheta}\rVert_{h}>0$
implies that $T$ is past-directed. ∎
Lemma (Lemma 2.7 (restated)).
Let $\tau$ be a time function on $M$. If $\tau$ is differentiable at a point $p\in M$, then the gradient vector $(\nabla\tau)_{p}$ is either past-directed causal or zero.
Proof.
We first claim that for every future-directed timelike vector $T\in T_{p}M$, $g((\nabla\tau)_{p},T)\geq 0$. Indeed, let $\gamma\colon(-\delta,\delta)\to M$ be a future-directed timelike curve with $\gamma(0)=p$ and $\gamma^{\prime}(0)=T$. Since $\tau$ is a time function, the composition $\tau\circ\gamma$ is strictly increasing. Hence, its derivative at zero satisfies $(\tau\circ\gamma)^{\prime}(0)\geq 0$. This derivative equals $\mathrm{d}_{p}\tau(T)=g((\nabla\tau)_{p},T)$, so the claim follows.
Suppose, for contradiction, that $(\nabla\tau)_{p}$ is spacelike, that is, $g((\nabla\tau)_{p},(\nabla\tau)_{p})>0$. Choose a future-directed timelike vector $T\in T_{p}M$ such that $g(T,(\nabla\tau)_{p})=0$, and define
$T^{\prime}=T+\lambda(\nabla\tau)_{p},$
where $\lambda<0$ is chosen so that
$g(T,T)+\lambda^{2}g((\nabla\tau)_{p},(\nabla\tau)_{p})<0.$
Let $\vartheta$ denote the time orientation. Then:
  1. (1)
    $T^{\prime}$ is timelike, since
    $g(T^{\prime},T^{\prime})$ $=g(T+\lambda(\nabla\tau)_{p},T+\lambda(\nabla\tau)_{p})$
    $=g(T,T)+2\lambda g(T,(\nabla\tau)_{p})+\lambda^{2}g((\nabla\tau)_{p},(\nabla\tau)_{p})$
    $=g(T,T)+\lambda^{2}g((\nabla\tau)_{p},(\nabla\tau)_{p})$
    $<0.$
  2. (2)
    $T^{\prime}$ is future-directed, since
    $g(T^{\prime},\vartheta)=g(T,\vartheta)+\lambda g((\nabla\tau)_{p},\vartheta)<0,$
    where we used that $g(T,\vartheta)<0$ because $T$ is future-directed, and that $g((\nabla\tau)_{p},\vartheta)\geq 0$ by the initial claim.
However,
$g((\nabla\tau)_{p},T^{\prime})$ $=g((\nabla\tau)_{p},T)+\lambda g((\nabla\tau)_{p},(\nabla\tau)_{p})$
$=\lambda g((\nabla\tau)_{p},(\nabla\tau)_{p})$
$<0,$
which contradicts the initial claim. Hence, $(\nabla\tau)_{p}$ cannot be spacelike and must be either causal or zero. Finally, if $(\nabla\tau)_{p}$ is causal, then it must be past-directed since $g((\nabla\tau)_{p},\vartheta)\geq 0$. ∎